### v2

# Pre-training Differentially Private Models with Limited Public Data

Abstract
The superior performance of large foundation models relies on the use of massive amounts of high-quality data, which often contain sensitive, private and copyrighted material that requires formal protection. While differential privacy (DP) is a prominent method to gauge the degree of security provided to the models, its application is commonly limited to the model fine-tuning stage, due to the performance degradation when DP is applied during the pre-training stage. Consequently, DP is yet not capable of protecting a substantial portion of the data used during the initial pre-training process. In this work, we provide a theoretical understanding of the efficacy of DP training by analyzing the per-iteration loss improvement, through the lens of Hessian matrix for large neural networks. We make a key observation that DP optimizers’ performance degradation can be significantly mitigated by the use of limited public data, which leads to a novel DP continual pre-training strategy. Empirically, using only 10% of public data, our strategy can achieve DP accuracy of 41.5% on ImageNet-21k (with $\epsilon=8$), as well as non-DP accuracy of 55.7% and 60.0% on downstream tasks Places365 and iNaturalist-2021, respectively, on par with state-of-the-art standard pre-training and substantially outperforming existing DP pre-trained models. Our DP pre-trained models are released in fastDP library (https://github.com/awslabs/fast-differential-privacy/releases/tag/v2.1).
1 Introduction
Large pre-trained models have been the backbone of computer vision and natural language processing. Finetuning or zero/few-shots learning (including in-context learning) based on these models can achieve superior performance. In particular, differentially private (DP) fine-tuning, such as full-parameter training, LoRA, Adapter, BiTFiT, and linear probing [89, 49, 14, 12, 58, 24], has shown to be almost as accurate as the standard non-DP fine-tuning on GPT [68, 6], ViT [27], and ResNet [39] models, while protecting the privacy of fine-tuning data. To be more specific, DP language models (GPT/BERT/RNN) and vision models are highly effective in defending against canary insertion attacks [66, 40, 42] and membership inference attacks [20, 69, 32, 16]; DP fine-tuned GPT2 also reduces the personally identifiable information leakage by $5\sim 10$ times compared with its non-DP counterpart [53].
These DP fine-tuned models all follow a two-stage training procedure, in which the first stage trains a model on large public datasets from scratch (without privacy protection), and the second stage fine-tunes on relatively small private datasets. However, a growing concern has been raised against the pre-training on the vast amount of web-collected data [18, 17, 79, 41, 62, 67]. The pre-trained models could memorize and re-generate the sensitive information in the training data, which include copyright content in books and codes, images of faces and nudity, and other personally identifiable information such as address and phone numbers, even if the data have been pre-processed by some content filters. While it is common to use close-source or proprietary datasets (e.g. JFT [75, 24, 58]) instead and to not release the models trained on these datasets, this approach renders the models not reproducible and may still violate the data privacy in a more implicit way.
Consequently, because of the uncertainty in seemingly safe-to-use public datasets, it is important to apply DP to the pre-training phase, which is computationally feasible, since DP optimization can be as efficient as the standard non-DP training [13, 38]. However, DP optimization without any public data suffers from slow convergence, and suboptimal performance: even on CIFAR10, the accuracy drops from $>95\%$ to $<70\%$ at $\epsilon=8$ [63]; on GPT2, the non-DP BLEU score degrades from 65.73 (quality often better than human) to 15.457 (hard to get the gist) at $\epsilon=3$ [49]. In short, pre-training and fine-tuning differ significantly in the amount of data and computation resources used, as well as in optimization setups. We summarize the difference in Table 1.
Table 1: Comparing DP pre-training and DP fine-tuning.
pre-training fine-tuning
Dataset size large small
Training iterations large small
(amount of compute)
Trainable model parameters 100% $0.1\sim 100$%
Major cause of DP noising DP clipping
performance degradation
1.1 Contributions
Our main contributions are listed below. We emphasize that we have taken the privacy accounting and the convergence (i.e. the dependence of loss on $\sigma(B,T,\epsilon),B$) jointly into consideration.
(1) We provide a new perspective of analyzing the loss improvement in DP training with different optimizers. Specifically, we propose a Hessian-based framework to analyze the per-iteration improvement on the generalization loss and the effects of per-sample gradient clipping, noising, and hyperparameter choices (e.g., batch size, learning rate).
(2) Based on the framework, we can analyze the effect that DP mechanisms, especially the DP noise but less so the DP clipping, have on pre-training and fine-tuning stages, leading to some theoretical justifications about why pre-training is more vulnerable to DP noise; see Figure 1 for some empirical comparisons, and Section 3 for some in-depth discussion.
(3) Our analysis suggests that the deceleration due to DP mechanisms can be mitigated by using a certain amount of public data. We then propose a DP continual pre-training strategy that continues the pre-training privately from a non-DP initialization. Our strategy is easily implementable, automatic, and effective, which obtains accuracy on upstream and downstream tasks that is on par with its non-DP variants, while significantly outperforming state-of-the-art DP pretraining algorithms (see Figure 2).
(a) Pre-training ViT-Base (CIFAR10,$\epsilon=1$)
(b) Fine-tuning ViT-Large (CIFAR100, $\epsilon=1$)
(c) Pre-training GPT2-Large (CodeParrot, $\epsilon=8$)
(d) Fine-tuning GPT2-Large (E2E, $\epsilon=1$)
Figure 1: Comparison among the convergence of standard SGD, clipped SGD without noise, noisy SGD without clipping, and DP-SGD in different tasks and training stages.
Figure 2: Summary of results in Section 5. First three figures compare the downstream and few-shot performance and the data efficiency (circle’s radius proportional to pre-training data size) of the DP pre-trained models; the last figure shows the performance of DP pre-trained models defending against privacy attacks (lower is stronger in defense).
1.2 Related works
This work is closely related to other works in DP convergence analysis, DP fine-tuning of large models, the continual training and the Hessian-based analysis. We refer to Appendix C for an extended discussion.
1.3 Notions & settings
[…]
We study DP optimization under a fixed computation budget (with a fixed number of samples), a fixed privacy budget (with a fixed DP guarantee), and limited availability of public data.
1.4 Computation budget
We consider a fixed computation budget, so that the training ends when a fixed number of samples $S:=BT$ have been processed. As $B$ increases, the number of iterations $T$ decreases linearly, while the per-iteration training time increases almost linearly. This is because foundation models are generally trained with large batch size, which requires the use of distributed learning and gradient accumulation [^1] The mini-batch of size $B$ (a.k.a. the logical batch size) is divided into $B/b$ micro-batches, where the micro-batch size is $b\ll B$ (a.k.a. the physical or per-GPU batch size, which determines the training speed and memory cost). When $b$ is fixed subject to GPU memory, the per-iteration training time is proportional to the number of micro-batches and thus to $B$.. For example, Vision Transformers (ViT, [27, 74] is trained with $B=4\textup{k}$, GPT-3  [6] and LLaMA [77] with $B\approx 2\textup{k}$, DP-ResNet/ViT with $B=4\textup{k}$ in [24] and $B=1\textup{M}$ in [58], DP-RoBERTa with $B=2\textup{k}$ in [89], and DP-GPT2 with $B=1\textup{k}$ in [49].
As a consequence, the batch size $B$ has nearly zero influence on the total training time, leaving its effect only on the convergence speed.
1.5 Privacy budget
Definition 1.1 ([29, 26]).
A randomized algorithm ${\mathcal{M}}$ is $(\varepsilon,\delta)$-DP if, for any two neighboring datasets ${\mathcal{S}},{\mathcal{S}}^{\prime}$ that differ by one sample and for any event ${\mathcal{E}}$,
${\mathbb{P}}[{\mathcal{M}}({\mathcal{S}})\in{\mathcal{E}}]\leqslant{\textnormal{e}}^{\varepsilon}{\mathbb{P}}\left[{\mathcal{M}}\left({\mathcal{S}}^{\prime}\right)\in{\mathcal{E}}\right]+\delta,\text{where }\delta<1/n.$
Figure 3: Noise levels by privacy accountants.
Figure 4: Per-sample gradient clipping in (3).
We consider a fixed privacy budget $(\epsilon,\delta)$, to be realized through the DP optimizers in (2). Essentially, a noise $\sigma\mathcal{N}(0,{\mathbf{I}})$ is injected to the gradient at each iteration, whereas the noise magnitude $\sigma(B)$ is determined by privacy accountants such as RDP (default) [59], PRV [34] and GDP [26, 8] (see Figure 4).
2 Understanding DP training through the lens of Hessian
In this section, we derive the impact of different components in DP training, including the per-sample clipping, the noising, and hyper-parameter choices. To do so, we characterize the per-iteration per-sample loss improvement through the lens of Hessian.
We aim to optimize the generalization loss (equivalently the expected risk) for any differentiable loss function, e.g. the cross entropy, mean squared error, or adversarial loss [33]:
$\min_{{\mathbf{w}}\in\mathbb{R}^{d}}L({\mathbf{w}})=\mathbb{E}_{x}[L({\mathbf{w}},x)].$ (1)
We denote the per-sample gradient ${\mathbf{g}}_{i}({\mathbf{w}}):=\frac{\partial L({\mathbf{w}},x_{i})}{\partial{\mathbf{w}}}\in\mathbb{R}^{d}$, the oracle gradient ${\mathbf{G}}({\mathbf{w}}):=\frac{\partial\mathbb{E}_{x}[L({\mathbf{w}},x)]}{\partial{\mathbf{w}}}\in\mathbb{R}^{d}$, and the oracle Hessian matrix ${\mathbf{H}}({\mathbf{w}}):=\frac{\partial^{2}\mathbb{E}_{x}[L({\mathbf{w}},x)]}{\partial{\mathbf{w}}^{2}}\in\mathbb{R}^{d\times d}$. In 2.1, samples follow identically and independently (i.i.d.) from some data distribution, with no restriction on the covariance structure $\mathbf{\Sigma}({\mathbf{w}})$.
Assumption 2.1.
Per-sample gradients ${\mathbf{g}}_{i}({\mathbf{w}})$ are i.i.d with
$\mathbb{E}[{\mathbf{g}}_{i}({\mathbf{w}})]={\mathbf{G}}({\mathbf{w}}),\quad\textup{Cov}({\mathbf{g}}_{i}({\mathbf{w}}))=\mathbf{\Sigma}({\mathbf{w}}).$
To simplify the notation, we drop the dependence on ${\mathbf{w}}$.
Consider the general DP optimizers, such as SGD and Adam [46], which update the parameters with the privatized gradient,
${\mathbf{g}}=\frac{\sum_{i\in{\mathcal{B}}}C_{i}{\mathbf{g}}_{i}+\sigma\mathcal{N}(0,{\mathbf{I}}_{d})}{B}\approx\frac{c\sum_{i\in{\mathcal{B}}}{\mathbf{g}}_{i}+\sigma\mathcal{N}(0,{\mathbf{I}}_{d})}{B}.$ (2)
Here $C_{i}:=C(\|{\mathbf{g}}_{i}\|;R)$ is the per-sample clipping factor that restricts the sensitivity of $\sum_{i}C_{i}{\mathbf{g}}_{i}$ to some constant $R$, i.e. $\|C_{i}{\mathbf{g}}_{i}\|\leq R$. We set $\|C_{i}{\mathbf{g}}_{i}\|\leq 1$ (thus omitting $R$ throughout the paper) following the re-parameterized gradient clipping [24] and the automatic (AUTO) clipping [14], whose $C_{i}$’s are listed below:
$C_{i,\text{re-param}}=\min\left\{\frac{1}{\|{\mathbf{g}}_{i}\|},\frac{1}{R}\right\}\text{ or }C_{i,\text{AUTO}}=\frac{1}{\|{\mathbf{g}}_{i}\|}.$ (3)
Figure 4 illustrates the values of $C_{i}$ under different clipping functions. Note that in (2), we employ a crucial approximation $c\approx\mathbb{E}[C_{i}]$ so as to unify the formula of DP and non-DP SGD in Remark 2.2. This approximation only holds when the directions of vectors $\sum_{i}C_{i}{\mathbf{g}}_{i}$ and $\sum_{i}{\mathbf{g}}_{i}$ are very close, i.e., there is little per-sample clipping bias. Such approximation is empirically validated in Figure 1, where we observe from the ‘SGD’ and ‘SGD+clipping’ curves that the convergence (without noising) is not much influenced by the bias of per-sample gradient clipping.
Remark 2.2.
Setting $c=1$ and $\sigma=0$, the gradient (2) reduces to the standard mini-batch gradient. Hence, the difference between SGD and DP-SGD is characterized by $(\sigma,c)$.
2.1 Per-iteration improvement of DP-SGD
Next, we characterize and analyze the per-iteration improvement of DP-SGD through the lens of Hessian, under different choices of hyperparameters and clipping functions: ${\mathbf{w}}_{t+1}={\mathbf{w}}_{t}-\eta{\mathbf{g}}_{t}$. The extension of the analysis to more general optimizers (e.g., DP-Adam) is given in Section 3.4. We are interested in minimizing the second-order Taylor approximation of $L({\mathbf{w}}-\eta{\mathbf{g}})$, which is sufficiently accurate since parameter updates are often quite small [57]. The loss improvement in one iteration can be approximated as:
$L({\mathbf{w}})-L({\mathbf{w}}-\eta{\mathbf{g}})\approx\eta{\mathbf{G}}^{\top}{\mathbf{g}}-\frac{\eta^{2}}{2}{\mathbf{g}}^{\top}{\mathbf{H}}{\mathbf{g}}.$
Taking the expectation of the right-hand side, we obtain the expected per-iteration loss improvement (derived in Section A.1):
$\Delta L$ $:=\eta{\mathbf{G}}^{\top}\mathbb{E}[{\mathbf{g}}]-\frac{\eta^{2}}{2}(\textup{tr}\left({\mathbf{H}}\textup{Cov}({\mathbf{g}})\right)+\mathbb{E}[{\mathbf{g}}]^{\top}{\mathbf{H}}\mathbb{E}[{\mathbf{g}}]).$ (4)
By applying Assumption 2.1 and (2), we have $\mathbb{E}[{\mathbf{g}}]=c{\mathbf{G}},\quad\textup{Cov}({\mathbf{g}})=c^{2}\mathbf{\Sigma}/B+\sigma^{2}/B^{2}.$ Substitute to (4), we obtain a quadratic function of $\eta$:
$\Delta L_{\textup{priv}}(\eta,B):=\eta c{\mathbf{G}}^{\top}{\mathbf{G}}-\frac{\eta^{2}}{2}\left(c^{2}{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}+\frac{c^{2}\textup{tr}({\mathbf{H}}\mathbf{\Sigma})}{B}+\frac{\sigma^{2}\textup{tr}({\mathbf{H}})}{B^{2}}\right).$ (5)
We denote the batch size used for SGD and DP-SGD as $B_{\text{non-DP}}$ and $B_{\text{DP}}$, respectively. Then by optimizing the learning rate $\eta$ 222 The optimal learning rate of DP-SGD cannot be used in practice because (i) the oracle ${\mathbf{G}}$ and ${\mathbf{H}}$ are unknown; (ii) it is data-dependent and hence violates the DP guarantee. Such an optimal learning rate is only used to help us understand the best possible per-iteration performance. , the per-sample and per-iteration improvement simplifies to
$\max_{\eta}\Delta L_{\textup{priv}}/B_{\text{DP}}=\Delta L^{\star}_{\textup{priv}}(B_{\text{DP}}):=\frac{1}{2}\frac{|{\mathbf{G}}|^{4}}{B_{\text{DP}}{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}+\textup{tr}({\mathbf{H}}\mathbf{\Sigma})+\sigma^{2}\textup{tr}({\mathbf{H}})/(B_{\text{DP}}c^{2})}.$ (6)
Given that the total number of processed samples is fixed at $S=BT$, (6) can be used as a metric to evaluate the data efficiency of DP and non-DP training with different $T,B$ and $(\epsilon,\delta)$.
2.2 Per-iteration improvement of vanilla SGD
We can analyze the loss improvement of standard SGD (which uses the full public data) as a sub-case of DP-SGD by substituting $c=1,\sigma=0$ into (5), according to Remark 2.2:
$\Delta L_{\textup{pub}}$ $:=\eta{\mathbf{G}}^{\top}{\mathbf{G}}-\frac{\eta^{2}}{2}\left(\frac{\textup{tr}({\mathbf{H}}\mathbf{\Sigma})}{B_{\text{non-DP}}}+{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}\right),\Delta L^{\star}_{\textup{pub}}(B):=\frac{1}{2}\frac{|{\mathbf{G}}|^{4}}{B_{\text{non-DP}}{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}+\textup{tr}({\mathbf{H}}\mathbf{\Sigma})}$ (7)
We visualize (6), (7), and their individual terms in Figure 5.
(a) DP pre-training
(b) DP fine-tuning
Figure 5: Illustration of different terms in (6) and (7). Left sub-plots depict the denominators in (6) and (7). Right sub-plots depict the whole terms and optimal batch sizes.
Implication 2.3 (Better DP mechanism helps).
From (6), it is clear that smaller $\sigma$ and larger $c$ (hence larger $C_{i}$) can help DP training. To reduce $\sigma$, we refer to Section 3.1 for a discussion of methods. For the clipping, under the same sensitivity bound, AUTO clipping [14] gives the largest $C_{i}$ among all clipping functions (see Figure 4), and therefore is more preferred to use in practice.
Implication 2.4 (Batch size should be large, but not too large).
As visualized by the red solid curves in Figure 5, there exists an optimal batch size (marked in red dashed vertical lines)
$B^{\star}_{\text{DP}}:=\text{argmax}_{B}\Delta L^{\star}_{\textup{priv}}(B)\approx\sqrt{\frac{\sigma^{2}\textup{tr}({\mathbf{H}})}{c^{2}{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}}}.$
Compared to previous DP literature, which encourages the batch size to be as a large as possible, our derivation of $B^{\star}_{\text{DP}}$ indicates a sweet pot: while we also support the use of large batch size, we highlight the data inefficiency if $B_{\text{DP}}$ is too large, a case that is often overlooked.
3 Impact of per-sample clipping and noising
In this section, we examine the effects of per-sample gradient clipping and noising on the DP training, leveraging the per-iteration per-sample loss improvement derived in Section 2. Specifically, we define and analyze a “decelerator” term that characterizes the slowdown by DP optimizers.
Comparing DP-SGD to SGD, we can attribute the slow convergence of DP optimization to the term $\sigma^{2}\textup{tr}({\mathbf{H}})/(Bc^{2})$ in (6), which is not present in the standard training (7). We refer to such a term as
[…]
which couples the effects of per-sample gradient clipping and noise through the trace of Hessian. We note that $\textup{tr}({\mathbf{H}})$ in (8) characterizes the curvature (i.e. sharpness or flatness) of the loss landscape, which strongly correlates with the downstream performance [51, 45, 96, 31].
Next, we discuss how the decelerator impacts the non-DP training, DP pre-training and fine-tuning.
3.1 No noise, (almost) no deceleration
When $\sigma=0$ (i.e., no DP noise), the decelerator vanishes and hence (6) reduces to (7), even if the per-sample gradient clipping is used. We empirically verified this in Figure 1 (see blue and black curves), where we see that the difference in convergence with or without clipping is negligible.
Given that DP noise is critical to the convergence speed, we highlight some techniques to reduce $\sigma$ under the same budget of $(\epsilon,\delta)$: the advances in privacy accounting theory can justify smaller noise; algorithms such as LoRA, low-pass and Kalman filters [93, 92] can reduce the effective noise.
3.2 DP pre-training can be vulnerable to noise
When $\sigma\neq 0$, the decelerator is non-zero. Therefore, DP training is slowed down by the noise; in Figure 1, SGD with noise (yellow and red curves) has worse performance than SGD without noise (black and blue curves). Furthermore, in pre-training, the decelerator is relatively large in the denominator of (6), i.e., $B{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}+\textup{tr}({\mathbf{H}}\mathbf{\Sigma})\leq\frac{\sigma^{2}\textup{tr}({\mathbf{H}})}{Bc^{2}}$ when $B\leq B^{*}_{\text{DP}}$ (see left sub-plot of Figure 5(a)), and therefore the slowdown can be significant.
Note that the deceleration issue cannot be resolved by increasing $B$. Although increasing $B$ improves the relative speed of DP convergence in comparison to non-DP, i.e., the decelerator decreases, it hurts the absolute speed since $B{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}$ increases, and thus the loss improvement (6) also worsens (c.f. right sub-plot of Figure 5(a)). Therefore, to design an efficient DP pre-training strategy, we must keep $B$ moderate and reduce the decelerator simultaneously.
3.3 DP fine-tuning is robust to noise
Empirical evidence has shown that DP fine-tuning is comparable to (though slightly worse than) the standard non-DP fine-tuning [89, 49, 24, 58, 10, 14, 12], despite that $\sigma\neq 0$. Such a phenomenon implies that comparing public and DP finetuning, we have $\Delta L^{\star}_{\textup{pub}}(B)\approx\Delta L^{\star}_{\textup{priv}}(B)$. That is, the decelerator becomes small after the public pre-training. This is conceptually illustrated in Figure 5(b), where the DP curve is close to the non-DP curve at moderate $B$ during fine-tuning, but not so during pre-training.
To understand the stark contrast between DP fine-tuning and DP pre-training, we plug in the optimal $B^{\star}_{\text{DP}}=\sqrt{\frac{\sigma^{2}\textup{tr}({\mathbf{H}})}{c^{2}{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}}}$ from 2.4 to $\Delta L^{\star}_{\textup{priv}}(B)$. Then, we have the optimal improvement of DP-SGD as
$\Delta L^{\star}_{\textup{priv}}(B^{\star}_{\text{DP}})=\frac{1}{2}\frac{|{\mathbf{G}}|^{4}}{2\sqrt{{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}\cdot\sigma^{2}\textup{tr}({\mathbf{H}})/c^{2}}+\textup{tr}({\mathbf{H}}\mathbf{\Sigma})}.$
Implication 3.1.
Notice that by choosing $B_{\text{non-DP}}=2B^{\star}_{\text{DP}}$ in (7), we have that DP-SGD with the optimal batch size is as fast as the standard SGD with twice the batch size,
$\Delta L^{\star}_{\textup{pub}}(2B^{\star}_{\text{DP}})=\Delta L^{\star}_{\textup{priv}}(B^{\star}_{\text{DP}}).$ (9)
Moreover, if $B^{\star}_{\text{DP}}$ is moderate, then DP-SGD can converge at similar speed to a fast converging SGD that uses a moderate batch size.
Remark 3.2.
From (7), we observe that non-DP training is data-efficient only if $B_{\text{non-DP}}{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}\ll\textup{tr}({\mathbf{H}}\mathbf{\Sigma})$. Otherwise, we can decrease $B_{\text{non-DP}}$ to effectively improve $\Delta L^{\star}_{\textup{pub}}$. Therefore, DP training is data-efficient only if $B^{\star}_{\text{DP}}\approx\frac{1}{2}B_{\text{non-DP}}\ll\frac{\textup{tr}({\mathbf{H}}\mathbf{\Sigma})}{2{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}}$, which holds in fine-tuning but not in early stage of pre-training [^3] An intuitive explanation is that $\frac{\textup{tr}({\mathbf{H}}\mathbf{\Sigma})}{{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}}=\frac{\mathbb{E}[({\mathbf{g}}_{i}-{\mathbf{G}})^{\top}{\mathbf{H}}({\mathbf{g}}_{i}-{\mathbf{G}})]}{{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}}=\frac{\mathbb{E}({\mathbf{g}}_{i}^{\top}{\mathbf{H}}{\mathbf{g}}_{i})}{\mathbb{E}({\mathbf{g}}_{i})^{\top}{\mathbf{H}}\mathbb{E}({\mathbf{g}}_{i})}-1$ resembles the variance of per-sample ${\mathbf{g}}_{i}$ in the space of ${\mathbf{H}}$, which decreases as the model learns the common representation. We illustrate the magnitude of the three terms in (6) for pre-training and fine-tuning stages in Figure 6.
Figure 6: Evolution of terms in (6) and (7) that explains the deceleration of DP optimization, during pre-training (left two) and fine-tuning (right two) ViT-Base on CIFAR100.
In the fine-tuning phase of Figure 6, $\textup{tr}({\mathbf{H}})$ quickly decreases and so does the decelerator. Hence a moderate $B_{\text{DP}}\approx 100$ can allow fast convergence. However, in the pre-training phase, $\textup{tr}({\mathbf{H}})$ increases to a large value within 5 epochs and remains for a long time (say epoch 5 to 40) before it decreases again. Consequently, DP convergence is initially fast but only for a short period and overall DP optimization is much slower than non-DP optimization, as shown in Figure 1(a)(c).
3.4 Extension to general optimizers
The analysis in the previous two sections can be easily extended to arbitrary optimizers as well as techniques such as weight decay, gradient clipping, and parameter-efficient fine-tuning (PEFT). Let $p$ be an optimizer’s post-processor of gradient and consider ${\mathbf{w}}_{t+1}={\mathbf{w}}_{t}-\eta p({\mathbf{g}}_{t}).$
For examples, following the notation in Pytorch library [65], we can write Adam and SGD with momentum ($\mu$) and weight decay ($\lambda$),
[…]
Specially, we consider optimizers such that $p(\cdot)$ is scale-invariant (i.e. $p(c{\mathbf{G}})=p({\mathbf{G}})$), such as SignSGD/Adam [4] or normalized SGD/LAMB [61, 55, 87]. Applying Assumption 2.1 and the optimal learning rate, we derive the expected per-sample per-iteration improvement in 3.3, leaving the details Section A.5.
Implication 3.3.
Suppose the post-processor $p(\cdot)$ is scale-invariant. Denote ${\mathbf{p}}=p({\mathbf{G}})$, ${\mathbf{p}}^{\prime}=p^{\prime}({\mathbf{G}})$, the per-sample per-iteration improvement $\max_{\eta}\Delta L(\eta)/B$ simplifies to
$\frac{1}{2}\frac{|{\mathbf{p}}^{\top}{\mathbf{G}}|^{2}}{B{\mathbf{p}}^{\top}{\mathbf{H}}{\mathbf{p}}+\textup{tr}({\mathbf{p}}^{\prime\top}{\mathbf{H}}{\mathbf{p}}^{\prime}\mathbf{\Sigma})+\sigma^{2}\textup{tr}({\mathbf{p}}^{\prime\top}{\mathbf{H}}{\mathbf{p}}^{\prime})/(Bc^{2})}$
Interestingly, similar to the decelerator of DP-SGD (8), the decelerator $\sigma^{2}\textup{tr}({\mathbf{p}}^{\prime\top}{\mathbf{H}}{\mathbf{p}}^{\prime})/(Bc^{2})$ of these DP optimizers also couples the per-sample gradient clipping, the noise and the Hessian, rendering the theoretical implications from DP-SGD extendable to general DP optimizers.
4 Continual pre-training with DP
4.1 Necessity of public data in DP pre-training
In this section, we propose the DP continual pre-training strategy and demonstrate that the deceleration by DP can be effectively mitigated by using a certain amount of public data. We consider the mixed data training that uses both public data (with subscript 0 for related hyperparameters) and private data (with subscript 1). Then SGD becomes
${\mathbf{g}}_{\alpha,t}:=\frac{\alpha_{t}}{B_{0}}\sum_{j\in{\mathcal{B}}_{0}}{\mathbf{g}}_{j,t}+\frac{(1-\alpha_{t})}{B_{1}}(\sum_{i\in{\mathcal{B}}_{1}}C_{i,t}{\mathbf{g}}_{i,t}+\sigma\mathcal{N}(0,{\mathbf{I}}_{d})).$
Here $\alpha_{t}\in[0,1]$ controls the ratio of privatized and non-privatized gradients, taking different forms by public training (OnlyPublic, $\alpha_{t}=1$), private training (OnlyPrivate, $\alpha_{t}=0$), DPMD [2], a tunable constant [32, 52], and Sample [30, 44].
Table 2: Summary of $\alpha_{t}$ by mixed data training methods.
Ours DPMD Sample OnlyPublic OnlyPrivate
$\mathbb{I}(t<sT)$ 1-$\cos\frac{\pi t}{2K}$ $\frac{n_{\textup{pub}}}{n_{\textup{pub}}+n_{\textup{priv}}}$ 1 0
Using the mixed gradient ${\mathbf{g}}_{\alpha}$ and following (4), we can show that expected loss improvement is a bivariate quadratic function of $(\eta,\alpha)$ (see Section A.3). After minimizing with respect to both variables, we obtain:
$\alpha^{\star}=\left(\frac{1}{c}\frac{\textup{tr}({\mathbf{H}}\mathbf{\Sigma})\cdot B_{1}/B_{0}}{\textup{tr}({\mathbf{H}}\mathbf{\Sigma})+\sigma^{2}\textup{tr}({\mathbf{H}})/(B_{1}c^{2})}+1\right)^{-1}$ (10)
Remark 4.1.
We see that the optimal choice in (10) gives $\alpha^{\star}\neq 0$, which indicates that the optimal improvement of mixed data training is larger than private training, i.e., $\Delta L^{\star}_{\text{OnlyPrivate}}<\Delta L^{\star}_{\textup{mixed}}$ and the public data helps. On the other hand, the fact that optimal $\alpha^{\star}\neq 1$ also indicates that $\Delta L^{\star}_{\text{OnlyPublic}}<\Delta L^{\star}_{\textup{mixed}}$ i.e., the private data helps.
4.2 DP Continual pre-training strategy overview
Motivated by Section 3 and Remark 4.1, we propose a two-phase DP continual pre-training strategy in Algorithm 1: public pre-training for $sT$ steps followed by private continual pre-training for $(1-s)T$ steps, which is equivalent to setting $\alpha_{t}=\mathbb{I}(t<sT)$ and the constant $10<s<$ 1 controls the portion of steps of public training.
Remark 4.2.
Although (10) suggests that $\alpha_{t}\in(0,1)$, we only set binary $\alpha_{t}\in\{0,1\}$ so as to not implement two data loaders and two back-propagation mechanisms simultaneously (DP v.s. standard). Note the methods in Table 2 are difficult to scale and not yet openly implemented on distributed systems due to memory and synchronization issues.
Figure 7: Convergence of GPT2-small on CodeParrot with different pre-training strategies ($\epsilon=8$).
We highlight that DP continual pre-training is as distinct from DP fine-tuning as their non-DP counter-parts, despite both methods extending the learning from a pre-trained phase: the continual pre-training learns common knowledge without adapting to a specific task and serves a richer foundation for many downstream tasks (see more discussion in Appendix C). In fact, we show in Section A.7 that the loss improvement of DP continual pre-training can be almost as fast as non-DP pre-training (FullyPublic), thus closing the utility gap.
The switching, i.e. selecting $s$, can be automatically determined by public statistics without using any DP budget. For example, we can use early stopping based on the loss or accuracy, i.e., we switch after the metrics on the validation dataset stop improving (see Figure 7). Alternatively, we can monitor $B^{\star}_{\text{DP}}$ and switch when it drops to a moderate value so that the DP training converges fast by (9).
5 DP vision foundation models on ImageNet
We leverage the DP continual pre-training strategy discussed in Section 4 to train vision transformers [27] on ImageNet datasets. We use ImageNet-1k (1.3M images, 1k classes; [25]) for public pre-training, then ImageNet-11k (formally known as ImageNet-21k-P [^4] ImageNet-11k improves the dataset quality and thus the resulting models (see Table 3 in [70]), by removing the infrequent classes from the full ImageNet-21k, thus remaining 78% of its 14M images and 48% of 21.8k classes., 11M images, 11k classes; [70]) for private pre-training. Notice that ImageNet-11k/21k is significantly harder to learn, with SOTA accuracy $\approx 47\%$ [70, 74] as compared to $>85\%$ for ImageNet-1k. We apply data augmentation including random flipping, contract shift, rotation, and resizing to $224\times 224$.
We evaluate our DP ViT on upstream and downstream tasks (including few-shot), achieving high accuracy under low computation budget (compared to existing DP pre-trained models in Table 3). In short, we show that unlocking more data (90% of full ImageNet), which cannot be used by DINO due to privacy concern, is significantly beneficial.
Table 3: Pre-training strategies of models. Standard non-DP training is marked in black; DP training is in green. $\dagger$ indicates self-supervised without using the labels. “Images $\times$” is the total number of images used (dataset size×$\times$×epochs). “Non-privacy” means no DP guarantee on a subset of training data due to the non-DP pre-training phase.
reference model pre-training continual training non-privacy images $\times$
TAN [71] ImageNet-1k 1.2B
(Syn)ViP [90] ViT-Base Shaders21k Shaders21k 1.3B
ViP [90] ViT-Base Shaders21k LAION400M Shaders21k 1.9B
DINO [19] ViT-Base ImageNet-1k ImageNet-1k 0.3B
Ours This work ViT-Base ImageNet-1k ImageNet-11k ImageNet-1k 0.7B
MIIL [70] ViT-Base ImageNet-1k ImageNet-11k ImageNet-11k 1.4B
Original [27] ViT-Base ImageNet-21k ImageNet-1k ImageNet-21k 1.2B
AugReg [74] ViT-Base ImageNet-1k ImageNet-21k ImageNet-21k 4.3B
NFnet-JFT [24] ResNet-50 JFT300M JFT300M 4.0B
5.1 Training strategy
For the public pre-training, we follow the self-supervised learning by [19] (self-distillation with no labels, or DINO), whereas the private continual pre-training is a supervised learning following [70] [^5] We observe that, using augreg (a supervised pre-training strategy) for public pre-training on ImageNet-1k and/or ImageNet-21k for private pre-training gives very similar result.. We employ AdamW optimizer with batch size $B=4096$ and learning rate $\eta=0.0002$ set by the line search. Our training strategy is similar to a concurrent work [90], with critical differences highlighted in Appendix C.
When automatically switching from public to private pre-training, the classification head (the last layer) is re-initialized because the number of classes is different in the two pre-training phases. This switching is triggered by early stopping. In the continual pre-training phase, we train with DP linear probing for 10 epochs then with DP full-parameter training for 20 epochs, with a total $\approx$100k training steps. This strategy achieves an upstream accuracy 41.5% on ImageNet-11k by our ViT-Base under $\epsilon=8$, and 39.8% under $\epsilon=2$.
5.2 Algorithm implementation
We employ fastDP library to apply the DP optimizers with automatic per-sample clipping [14] and layer-wise clipping style [9]. Specifically, the DP optimization is under the multi-GPU distributed system, using DP-ZeRO [7] and mixed-precision training. We calibrate the DP noise using the improved Renyi accountant.
5.3 Downstream performance
Our DP pre-training learns highly transferable representations, demonstrated through the strong performance on a list of downstream datasets. We summarize in Table 3 a number of models from previous literature for comparison. The most informative baselines are DINO and MIIL, since our DP models continue the pre-training from DINO, following the strategy of MIIL. Next, we mark DP pre-trained models in green and leave non-DP ones in black.
In Table 4, we compare different pre-training strategies (all non-DP except ours) leveraging the same dataset and same model architecture – ViT-Base (86M param). Our evaluation shows that DP pre-trained models achieve high downstream accuracy under standard and non-DP fine-tuning: 98.4% on CIFAR10, 90.2% on CIFAR100, 86.5% on Food101 and 96.8% on SVHN. Our DP continual pre-training clearly improves upon DINO, with $+0.3\sim 2.0\%$ on accuracy, and is comparable to the non-DP pre-trained MIIL that uses twice the data size (1.4B v.s. our 0.7B).
Table 4: Standard/DP fine-tuning accuracy with the same architecture (ViT-Base) and pre-training dataset (ImageNet-21k) up to subsampling and preprocessing. Number of processed images by each model is indicated in the parenthesis.
Pre-training Fine-tuning CIFAR10 CIFAR100 Food101 SVHN
non-DP $\epsilon=8$ $\epsilon=2$ non-DP $\epsilon=8$ $\epsilon=2$ non-DP $\epsilon=8$ $\epsilon=2$ non-DP $\epsilon=8$ $\epsilon=2$
DINO (0.3B) 98.1 97.2 97.0 88.2 84.7 82.7 85.2 77.2 73.5 96.2 91.7 90.3
Oursϵ=2 (0.7B) 97.8 96.6 96.1 88.8 83.1 81.1 84.8 75.5 72.5 96.3 91.3 90.1
Oursϵ=8 (0.7B) 98.4 97.2 96.9 90.2 85.0 82.8 86.5 78.4 75.3 96.8 92.5 91.3
MIIL (1.4B) 98.8 98.5 98.2 91.4 90.9 89.2 87.2 84.5 83.0 96.8 93.3 92.0
ViT_base (1.2B) 98.9 98.3 98.1 92.6 89.9 88.2 89.4 85.5 83.1 96.9 93.5 92.5
AugReg (4.3B) 98.9 98.8 98.5 93.1 91.2 90.4 90.2 87.6 85.7 96.9 93.8 92.5
In Table 6, when the downstream tasks (non-DP) are more challenging with only a few data samples to learn from, our DP model substantially outperforms previous DP pre-trained models across all settings, for example, by $+19\sim 38\%$ on CIFAR100 when compared to ViP and TAN. We attribute the success to the high quality of pre-training data, i.e. ImageNet-1k/11k, in contrast to Shaders (by comparing DINO to (Syn)ViP) and LAION (by comparing the improvement from DINO to ours and from (Syn)ViP to ViP).
Table 5: Few-shot accuracy of DP pre-trained models (TAN, ViP and ours) and their non-DP initialization.
Aircraft Aircraft CIFAR100 CIFAR100 fine-tune
(10-shot) (20-shot) (10-shot) (30-shot) epochs
TANϵ=8 22.84 37.93 27.78 42.35 200
(Syn)ViP 21.79 46.85 38.96 55.84 200
ViPϵ=8 31.62 53.05 40.95 57.52 200
DINO 32.04 45.61 47.31 66.92 100
Oursϵ=2 36.42 48.27 64.74 74.62 100
Oursϵ=8 42.57 57.15 65.26 76.38 100
Table 6: Linear-probing accuracy (non-DP) of pre-trained models, except “full” indicating full-parameter.
ImageNet-1k Places365 iNat2021 fine-tune
# images 1M 1.8M 2.7M epochs
TANϵ=8 49.0 40.5 31.7 90 / 90 / 90
(Syn)ViP 49.8 43.2 32.4 90 / 90 / 90
ViPϵ=8 55.7 46.1 38.1 90 / 90 / 90
DINO 76.1 52.1 43.5 8 / 5 / 10
Oursϵ=2 76.2 52.5 46.5 8 / 5 / 10
Oursϵ=8 77.9 53.0 49.1 8 / 5 / 10
Oursϵ=2(full) 78.0 55.6 57.2 8 / 5 / 10
Oursϵ=8(full) 78.5 55.7 60.0 8 / 5 / 10
NFnet-JFT 74.1 54.5 10 / 26 / —
In Table $6$For ImageNet-1k, we strike through some results because this dataset was used in the pre-training. It is only meaningful to compare our models to DINO and observe the benefit of continual pre-training, but not to others., we extend the evaluation of full fine-tuning and linear probing to SOTA non-DP baselines and to million-image scale Our DP model achieves 55.7% (+9.6% over ViP) on Places365 with 1.8M images and 60.0% (+21.9% over ViP) on iNat2021 with 2.7M images. The current non-DP SOTA is 57-60% on Places365 [24, 27] and about 64% on iNat2021 [80, 60] after pre-training on $2.7\sim 4\textup{B}$ images. This showcases the effectiveness of DP pre-training as our models only leverage 0.7B images.
5.4 Privacy protection
We employ a white-box membership inference attack (MIA) to evaluate the data protection by our DP pre-training: 1) for each image in ImageNet-11k, we compute its output logits and loss, which serves as the feature of the MIA dataset; 2) we randomly select $50\%$ of the testing images and the same number of training images ($522,496$ samples) as the MIA test set, and the rest data as MIA train set; 3) we label the training images as class “$1$” and testing images as class “00”. This creates the MIA dataset with 11k features and binary labels.
Table 7: Membership inference attack results. Values closer to $0.5$ indicate better privacy protection.
Accuracy Precision Recall F1 AUC
Oursϵ=2 50.1% 50.1% 47.5% 0.49 0.50
Oursϵ=8 $51.3$% $51.6$% $64.4$% $0.57$ $0.51$
Oursϵ=∞ $53.8$% $54.1$% $71.2$% $0.62$ $0.54$
MIIL $54.0$% $58.5$% $77.3$% $0.67$ $0.62$
We fit a logistic regression with the MIA training set to classify whether an image belongs to the training set of ImageNet-11k (class “$1$”) or not. We report the results on the MIA testing set in Table 7, showing the effectiveness of DP protection when $\epsilon\leq 8$.
6 Discussion
In this paper, we conduct an insightful and unified convergence analysis on DP optimization. Specifically, we identify the decelerator (8) of DP training as a result of the per-sample gradient clipping, the noise and the Hessian, which can be significantly mitigated by a small amount ($<10\%$) of public training. Consequently, we propose DP continual pre-training that is almost as accurate and implementable as the fully public pre-training.
Acknowledgement
The work of Xinwei Zhang was partially done while interning at Amazon. Mingyi Hong holds concurrent appointments as an Amazon Scholar and as a faculty at the University of Minnesota. This paper describes their work performed at Amazon.
Appendix A Derivation and proofs
A.1 Derivation of Equation 4
$\Delta L$ $:=\eta{\mathbf{G}}^{\top}\mathbb{E}[{\mathbf{g}}]-\frac{\eta^{2}}{2}\mathbb{E}[{\mathbf{g}}^{\top}{\mathbf{H}}{\mathbf{g}}]$
=$\stackrel{{\scriptstyle(a)}}{{=}}\eta{\mathbf{G}}^{\top}\mathbb{E}[{\mathbf{g}}]-\frac{\eta^{2}}{2}\textup{tr}\left({\mathbf{H}}\mathbb{E}[{\mathbf{g}}{\mathbf{g}}^{\top}]\right)$
=$\stackrel{{\scriptstyle(b)}}{{=}}\eta{\mathbf{G}}^{\top}\mathbb{E}[{\mathbf{g}}]-\frac{\eta^{2}}{2}\textup{tr}\left({\mathbf{H}}\textup{Cov}({\mathbf{g}})\right)+\textup{tr}\left({\mathbf{H}}\mathbb{E}[{\mathbf{g}}]\mathbb{E}[{\mathbf{g}}]^{\top}\right)$
=$\stackrel{{\scriptstyle(c)}}{{=}}\eta{\mathbf{G}}^{\top}\mathbb{E}[{\mathbf{g}}]-\frac{\eta^{2}}{2}(\textup{tr}\left({\mathbf{H}}\textup{Cov}({\mathbf{g}})\right)+\mathbb{E}[{\mathbf{g}}]^{\top}{\mathbf{H}}\mathbb{E}[{\mathbf{g}}])$
where $(a)$ uses equations (13) and then (12) to the second term; $(b)$ separates the expectation of ${\mathbf{g}}{\mathbf{g}}^{\top}$ into its mean squared and covariance and uses (11); $(c)$ uses equations (12) and then (13) to the last term.
Property of the trace:
[…]
A.2 Explanation for 2.4
We will leverage GDP to validate that $\sigma(B)^{2}/B\approx O(1/B)$, given that most of other privacy accountants are numerical and hard to interpret. We note that $\mu$-GDP has an one-to-one mapping with the $(\varepsilon,\delta)$-DP:
$\delta=\Phi(-\frac{\epsilon}{\mu}+\frac{\mu}{2})+\Phi(-\frac{\epsilon}{\mu}-{\textnormal{e}}^{\epsilon}\frac{\mu}{2})$
where $\Phi$ is the normal cumulative distribution function.
Lemma A.1.
Given an iterative algorithm with $\ell_{2}$ sensitivity $1$ at each iteration, which uniformly samples the data in dataset of size $n$ with ratio $\frac{B}{n}$, by injecting Gaussian noise ${\mathcal{N}}(0,\sigma^{2}{\mathbf{I}})$ to the output of the algorithm at each iteration, it satisfies $\mu$-GDP with
$\mu=\frac{B}{n}\sqrt{T({\textnormal{e}}^{1/\sigma(B)^{2}}-1)}=\sqrt{BS({\textnormal{e}}^{1/\sigma(B)^{2}}-1)}/n,$
where $S$ denotes the fixed computation budget.
The proof is equivalent to that in Section 2.4 in [26].
By Taylor expansion over $B$,
$\sigma(B)^{2}=\frac{1}{\log(\frac{\mu^{2}n^{2}}{BS}+1)}=\frac{BS}{\mu^{2}n^{2}}+\frac{1}{2}+O(\frac{1}{B})\longrightarrow\frac{\sigma(B)^{2}}{B}=\frac{S}{\mu^{2}n^{2}}+\frac{1}{2B}+o(\frac{1}{B})$
In the pre-training regime, $S/n$ is the number of epochs (usually between 1 and 300) and the sample size $n$ is huge, i.e. $n>10^{7}$ for ImageNet and $n>10^{1}2$ for large language model training. Hence the first term is negligible and $\sigma(B)^{2}/B\approx 0.5/B$.
A.3 Proof of results in Section 3.3
Proof.
Leveraging (2), we have
$\mathbb{E}({\mathbf{g}})$ $=\alpha{\mathbf{G}}+(1-\alpha)c{\mathbf{G}},\;\textup{Cov}({\mathbf{g}})$ $=\alpha^{2}\frac{\mathbf{\Sigma}}{B_{0}}+(1-\alpha)^{2}(\frac{c^{2}\mathbf{\Sigma}}{B_{1}}+\frac{\sigma^{2}{\mathbf{I}}}{B_{1}^{2}})$
Substituting in the mixed gradient ${\mathbf{g}}_{\alpha}$ and following (4), we can write down the expected improvement of mixed data training:
$\Delta L_{\textup{mixed}}$ $=\eta(\alpha+(1-\alpha)c){\mathbf{G}}^{\top}{\mathbf{G}}-\frac{\eta^{2}(1-\alpha)^{2}\sigma^{2}}{2B_{1}^{2}}\textup{tr}({\mathbf{H}})$
$-\frac{\eta^{2}}{2}(\frac{\alpha^{2}}{B_{0}}+\frac{(1-\alpha)^{2}}{B_{1}}c^{2})\textup{tr}({\mathbf{H}}\mathbf{\Sigma})$
$-\frac{\eta^{2}}{2}(\alpha+(1-\alpha)c)^{2}{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}},$
which is a bivariate quadratic function $\Delta L_{\textup{mixed}}$ to be minimized over $(\eta,\alpha)$. To make the presentation easier, we denote $\eta_{0}=\eta\alpha,\eta_{1}=\eta(1-\alpha)$, and the above equation can be rewritten as:
$\Delta L_{\textup{mixed}}=(\eta_{1}c+\eta_{0}){\mathbf{G}}^{\top}{\mathbf{G}}-\frac{1}{2}\frac{\eta_{1}^{2}\sigma^{2}}{B_{1}^{2}}\textup{tr}({\mathbf{H}})-\frac{1}{2}(\frac{\eta_{1}^{2}}{B_{1}}c^{2}+\frac{\eta_{0}^{2}}{B_{0}})\textup{tr}({\mathbf{H}}\mathbf{\Sigma})-\frac{1}{2}(\eta_{1}c+\eta_{0})^{2}{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}.$
Note that if we set $\eta_{1}=0$ or $\eta_{0}=0$, we essentially train the model with only public data or private data, respectively.
can be written as a bivariate quadratic function over $(\eta_{0},\eta_{1})$:
$\Delta L_{\textup{mixed}}=-(A\eta_{0}^{2}+B\eta_{1}^{2}+C\eta_{0}+D\eta_{1}+E\eta_{0}\eta_{1})$
in which
$A$ $=\frac{1}{2}{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}+\frac{1}{2}\frac{\textup{tr}({\mathbf{H}}\mathbf{\Sigma})}{B_{0}},$
$B$ $=\frac{1}{2}c^{2}{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}+\frac{1}{2}c^{2}\frac{\textup{tr}({\mathbf{H}}\mathbf{\Sigma})}{B_{1}}+\frac{1}{2}\frac{\sigma^{2}\textup{tr}({\mathbf{H}})}{B_{1}^{2}},$
$C$ $=-{\mathbf{G}}^{\top}{\mathbf{G}},$
$D$ $=-c{\mathbf{G}}^{\top}{\mathbf{G}},$
$E$ $=c{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}.$
The maximizer is
$\eta_{0}=-\frac{2BC-DE}{4AB-E^{2}},\eta_{1}=-\frac{2AD-CE}{4AB-E^{2}}.$
Hence
$\frac{1-\alpha^{*}}{\alpha^{*}}=\frac{2AD-CE}{2BC-DE}\Longrightarrow\alpha^{*}=1/\left(\frac{2AD-CE}{2BC-DE}+1\right)$
Finally
$\alpha^{*}=\left(\frac{c\textup{tr}({\mathbf{H}}\mathbf{\Sigma}){\mathbf{G}}^{\top}{\mathbf{G}}/B_{0}}{c^{2}\textup{tr}({\mathbf{H}}\mathbf{\Sigma}){\mathbf{G}}^{\top}{\mathbf{G}}/B_{1}+\sigma^{2}\textup{tr}({\mathbf{H}}){\mathbf{G}}^{\top}{\mathbf{G}}/B_{1}^{2}}+1\right)^{-1}=\left(\frac{1}{c}\frac{\textup{tr}({\mathbf{H}}\mathbf{\Sigma})\cdot B_{1}/B_{0}}{\textup{tr}({\mathbf{H}}\mathbf{\Sigma})+\sigma^{2}\textup{tr}({\mathbf{H}})/(B_{1}c^{2})}+1\right)^{-1}$
A.4 Explanation of Remark 3.2
From (6), the optimal batch size for DP-SGD is
$\operatorname*{arg\,max}_{B}\frac{1}{2}\frac{|{\mathbf{G}}|^{4}}{B{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}+\textup{tr}({\mathbf{H}}\mathbf{\Sigma})+\sigma^{2}\textup{tr}({\mathbf{H}})/(Bc^{2})}$
$=\operatorname*{arg\,min}_{B}B{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}+\textup{tr}({\mathbf{H}}\mathbf{\Sigma})+\sigma^{2}\textup{tr}({\mathbf{H}})/(Bc^{2})\approx\sqrt{\frac{\sigma^{2}\textup{tr}({\mathbf{H}})}{c^{2}{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}}}$
which minimizes (6) to
[…]
where $B_{\text{DP}}^{*}:=\sqrt{\frac{\sigma^{2}\textup{tr}({\mathbf{H}})}{c^{2}{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}}}$.
Note this is equivalent to applying $B_{\text{non-DP}}:=2\sqrt{\frac{\sigma^{2}\textup{tr}({\mathbf{H}})}{c^{2}{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}}}=2B_{\text{DP}}^{*}$ on (7),
$\text{SGD}(B_{\text{non-DP}})\equiv\text{SGD}(2B_{\text{DP}}^{*})\equiv\text{DP-SGD}(B_{\text{DP}}^{*}).$
We emphasize that generally
$\text{SGD}(2B_{\text{DP}})\not\equiv\text{DP-SGD}(B_{\text{DP}}).$
A.5 Loss improvement of general DP optimizers – 3.3
From ${\mathbf{w}}_{t+1}={\mathbf{w}}_{t}-\eta p({\mathbf{g}}_{t})$, the expected per-iteration loss improvement becomes
$\Delta L$ $=\eta{\mathbf{G}}^{\top}\mathbb{E}[p({\mathbf{g}})]-\frac{\eta^{2}}{2}(\textup{tr}\left({\mathbf{H}}\textup{Cov}(p({\mathbf{g}}))\right)+\mathbb{E}[p({\mathbf{g}})]^{\top}{\mathbf{H}}\mathbb{E}[p({\mathbf{g}})]).$ (14)
Applying Assumption 2.1 and delta method, with ${\mathbf{G}}=\frac{\partial L}{\partial{\mathbf{w}}}$, we have
$\sqrt{B}(p({\mathbf{g}})-p(c{\mathbf{G}}))=\sqrt{B}p^{\prime}(c{\mathbf{G}})\cdot({\mathbf{g}}-c{\mathbf{G}})+o_{p}(1)\to\mathcal{N}\left(0,p^{\prime}(c{\mathbf{G}})[c^{2}\mathbf{\Sigma}+\sigma^{2}/B]p^{\prime}(c{\mathbf{G}})^{\top}\right)$
where $p^{\prime}(c{\mathbf{G}})\equiv\frac{\partial p(c{\mathbf{G}})}{\partial c{\mathbf{G}}}\in\mathbb{R}^{d\times d}$. Hence
$\mathbb{E}(p({\mathbf{g}}))\approx p(c{\mathbf{G}}),\textup{Cov}(p({\mathbf{g}}))\approx p^{\prime}(c{\mathbf{G}})\cdot\left(c^{2}\mathbf{\Sigma}/B+\sigma^{2}/B^{2}\right)\cdot p^{\prime}(c{\mathbf{G}})^{\top}$
Next, the expected improvement contributed by one sample is a quadratic function of $\eta$:
$\Delta L/B\approx$ $\left(\eta p(c{\mathbf{G}})^{\top}{\mathbf{G}}-\frac{\eta^{2}}{2}\frac{\sigma^{2}\textup{tr}[p^{\prime}(c{\mathbf{G}})^{\top}{\mathbf{H}}p^{\prime}(c{\mathbf{G}})]}{B^{2}}\right.$
$\qquad\left.-\frac{\eta^{2}}{2}\frac{c^{2}\textup{tr}[p^{\prime}(c{\mathbf{G}})^{\top}{\mathbf{H}}p^{\prime}(c{\mathbf{G}})\mathbf{\Sigma}]}{B}-\frac{\eta^{2}}{2}p(c{\mathbf{G}})^{\top}{\mathbf{H}}p(c{\mathbf{G}})\right)/B$
Applying the optimal learning rate, the optimal per-sample loss improvement $\Delta L/B$ at each iteration simplifies to
$\frac{1}{2}\frac{|p(c{\mathbf{G}})^{\top}{\mathbf{G}}|^{2}}{Bp(c{\mathbf{G}})^{\top}{\mathbf{H}}p(c{\mathbf{G}})+c^{2}\textup{tr}(p^{\prime}(c{\mathbf{G}})^{\top}{\mathbf{H}}p^{\prime}(c{\mathbf{G}})\mathbf{\Sigma})+\sigma^{2}\textup{tr}(p^{\prime}(c{\mathbf{G}})^{\top}{\mathbf{H}}p^{\prime}(c{\mathbf{G}}))/B}$
In the special case that $p$ is scale-invariant, e.g. in adaptive optimizers like Adam or in SignSGD, we get $p^{\prime}(c{\mathbf{G}})=\frac{\partial p(c{\mathbf{G}})}{\partial c{\mathbf{G}}}=\frac{\partial p({\mathbf{G}})}{\partial{\mathbf{G}}}\frac{\partial{\mathbf{G}}}{\partial c{\mathbf{G}}}=p^{\prime}({\mathbf{G}})/c$ and thus
$\frac{1}{2}\frac{|p({\mathbf{G}})^{\top}{\mathbf{G}}|^{2}}{Bp({\mathbf{G}})^{\top}{\mathbf{H}}p({\mathbf{G}})+\textup{tr}(p^{\prime}({\mathbf{G}})^{\top}{\mathbf{H}}p^{\prime}({\mathbf{G}})\mathbf{\Sigma})+\sigma^{2}\textup{tr}(p^{\prime}({\mathbf{G}})^{\top}{\mathbf{H}}p^{\prime}({\mathbf{G}}))/(Bc^{2})}$
A.6 Improvement of performance measures other than the optimization loss
We now consider two extended cases when the optimization loss is different to the performance measures. For example, the model may be trained via the cross-entropy loss but measured on 0-1 accuracy.
optimization performance example
$L$ $L$ vanilla $L$ is cross-entropy
$L$ $L_{\text{other}}$ vanilla $L$ is cross-entropy; $L_{\text{other}}$ is BLEU or accuracy
$L_{\text{mod}}$ $L$ adversarial training $L_{\text{mod}}$ is adversarial loss; $L$ is cross-entropy
Table 8: DP/non-DP optimization when the optimization loss is different to the performance measures.
For the first case, we analyze many performance measures of DP models (denoted as $L_{\textup{other}}$) beyond the optimization loss $L$. For instance, foundation models trained on cross-entropy loss can be evaluated on classification accuracy, F1 score, BLEU [64], ROGUE [50], fairness [37], calibration [35], adversarial robustness, etc.
We demonstrate that our analysis in previous sections indeed generalizes: equivalent to (5), we have
$\Delta L_{\textup{other}}(\eta)=\eta{\mathbf{G}}_{\textup{other}}^{\top}\mathbb{E}[{\mathbf{g}}]-\frac{\eta^{2}}{2}\mathbb{E}[{\mathbf{g}}^{\top}{\mathbf{H}}_{\textup{other}}{\mathbf{g}}]$
where ${\mathbf{G}}_{\textup{other}},{\mathbf{H}}_{\textup{other}}$ are the oracle gradient and Hessian of $L_{\textup{other}}$. The per-sample per-iteration improvement, $\max_{\eta}\Delta L_{\textup{other}}/B$, simplifies to
$\frac{1}{2}\frac{|{\mathbf{G}}_{\textup{other}}^{\top}{\mathbf{G}}|^{2}}{B{\mathbf{G}}^{\top}{\mathbf{H}}_{\textup{other}}{\mathbf{G}}+\textup{tr}({\mathbf{H}}_{\textup{other}}\mathbf{\Sigma})+\sigma^{2}\textup{tr}({\mathbf{H}}_{\textup{other}})/(Bc^{2})},$
with the decelerator being $\sigma^{2}\textup{tr}({\mathbf{H}}_{\textup{other}})/(Bc^{2})$ in place of (8). As implied by Section 3.3, we expect the public pre-training also mitigates this decelerator, so that well-trained DP models and non-DP models should be equally performant. Empirically speaking, DP fine-tuning has shown to be as accurate [49, 24], adversarially robust [15], calibrated [11], and fair [5] as the standard fine-tuning.
For the second case, e.g. adversarial, sharpness-aware, fairness-aware or calibration-aware training, the optimization is on the modified loss $\frac{\partial L_{\text{mod}}}{\partial{\mathbf{w}}}$ but the performance is measured on the vanilla loss $\frac{\partial L}{\partial{\mathbf{w}}}$. For instance, we want an adversarially trained model to be sufficiently accurate. To be specific, FGSM and PGD ($L_{2}/L_{\infty}$ perturbation with norm $\rho$) lead to the modified loss
$L_{\text{mod}}:=\max_{||\xi||\leq\rho}\mathbb{E}_{x}[L({\mathbf{w}},x+\xi)],{\mathbf{G}}_{\text{mod}}=\frac{\partial L_{\text{mod}}}{\partial{\mathbf{w}}}.$
DP adversarial training applies on
${\mathbf{p}}=\frac{1}{B}(\sum_{i}C_{i}{\mathbf{g}}_{\text{mod},i}+\sigma\mathcal{N}(0,{\mathbf{I}}_{d}))\approx\frac{1}{B}(c\sum_{i}{\mathbf{g}}_{\text{mod},i}+\sigma\mathcal{N}(0,{\mathbf{I}}_{d}))$
where per-sample gradient is ${\mathbf{g}}_{\text{mod},i}:=\frac{\partial\max_{||\xi||\leq\rho}L({\mathbf{w}},x_{i}+\xi)}{\partial{\mathbf{w}}}$. Note that the modified gradient may be post-processed by optimizers like AdamW. To be clear, only in this section, we write ${\mathbf{p}}=\frac{\partial L_{\text{mod}}}{\partial{\mathbf{w}}}$ and omit the post-processing of the optimizer (i.e. we only show for SGD).
Next, the expected per-iteration loss improvement becomes
$\Delta L$ $=\eta{\mathbf{G}}^{\top}\mathbb{E}[{\mathbf{p}}]-\frac{\eta^{2}}{2}(\textup{tr}\left({\mathbf{H}}\textup{Cov}({\mathbf{p}})\right)+\mathbb{E}[{\mathbf{p}}]^{\top}{\mathbf{H}}\mathbb{E}[{\mathbf{p}}]).$ (15)
Applying Assumption 2.1, we have
$\mathbb{E}[{\mathbf{p}}]=c{\mathbf{G}}_{\text{mod}},\textup{Cov}({\mathbf{p}})=c^{2}\mathbf{\Sigma}_{\text{mod}}/B+\sigma^{2}/B^{2}$
Hence, the expected per-iteration improvement contributed by one sample is a quadratic function of $\eta$:
$\Delta L/B$ $:=(\eta c{\mathbf{G}}^{\top}{\mathbf{G}}_{\text{mod}}-\frac{\eta^{2}}{2}\frac{\sigma^{2}\textup{tr}({\mathbf{H}})}{B^{2}}-\frac{\eta^{2}}{2}\frac{c^{2}\textup{tr}({\mathbf{H}}\mathbf{\Sigma}_{\text{mod}})}{B}-\frac{\eta^{2}}{2}c^{2}{\mathbf{G}}_{\text{mod}}^{\top}{\mathbf{H}}{\mathbf{G}}_{\text{mod}})/B$
Applying the optimal learning rate, the optimal per-sample per-iteration improvement, $\max_{\eta}\Delta L/B$, simplifies to
$\frac{1}{2}\frac{|{\mathbf{G}}^{\top}{\mathbf{G}}_{\text{mod}}|^{2}}{B{\mathbf{G}}_{\text{mod}}^{\top}{\mathbf{H}}{\mathbf{G}}_{\text{mod}}+\textup{tr}({\mathbf{H}}\mathbf{\Sigma}_{\text{mod}})+\sigma^{2}\textup{tr}({\mathbf{H}})/(Bc^{2})}$ (16)
Again, the decelerator is the same as (8). Therefore, DP training will be as good as the standard training when the decelerator is small under the modified loss. This supports the observation that DP adversarial training can be accurate in [15], if DP natural training is comparable to non-DP natural training.
A.7 Explaining the effectiveness of DP continual pre-training
Taking a step further than Remark 4.1, we claim that DP continual pre-training can converge as fast as non-DP pre-training (FullyPublic), conditioning on that the non-DP initialization is sufficiently strong: running $sT$ iterations of public training followed by the DP training is only marginally weaker than FullyPublic in (7):
[…]
$\lessapprox\sum_{t=1}^{sT}\frac{|{\mathbf{G}}_{t}|^{4}}{{\mathbf{G}}^{\top}_{t}{\mathbf{H}}_{t}{\mathbf{G}}_{t}+\frac{\textup{tr}({\mathbf{H}}_{t}\mathbf{\Sigma}_{t})}{B}}+\sum_{t=sT}^{T}\frac{|{\mathbf{G}}_{t}|^{4}}{{\mathbf{G}}^{\top}_{t}{\mathbf{H}}_{t}{\mathbf{G}}_{t}+\frac{\textup{tr}({\mathbf{H}}_{t}\mathbf{\Sigma}_{t})}{B}},$
given that the decelerator (8) is small for $t>sT$.
Appendix B Experiment settings and additional details
In this section, we provide additional descriptions and experiment details for the numerical results in the main paper.
B.1 Datasets
We use the following datasets throughout this paper.
  • ImageNet: ImageNet-21k [25] is the full dataset with various releases which contain 14.2M images from 21,841 classes. ImageNet-1k is a subset of ImageNet-21k using 1k high-level classes. We use the ILSVRC2012 version, which contains 1.2 million training images and 100000 test images. ImageNet-11k [70] is another subset of ImageNet-21k that removes invalid/infrequent classes, retaining 11.1M images (train:test=10.5M :0.52M ) and 10,450 classes in the Winter21 version.
  • CIFAR-10/CIFAR-100: 50,000 training and 10,000 test images, with 10 or 100 classes, respectively.
  • Food101: it contains 101 classes of food, with 101,000 images in total, 1000 images per class (training 750 and test 250).
  • SVHN: it contains 10 classes of digits in natural scene images, with 73257 training and 26032 test images.
  • Aircraft: FGVCAircraft dataset contains 3334 training and $3$ test images with 100 classes.
  • Places365: Places365-Standard dataset contains 1.8 million training and 36000 test images from 365 scene classes.
  • iNat2021: iNaturalist 2021 dataset contains 10,000 classes of species, with 2.7 million training and 0.1 million test images.
  • CodeParrot: a GitHub dataset of about 180 GB containing roughly 20 million Python files. We use 3% of these files, that is 606,720 rows of training and 3322 rows of test data. We take sequence length 128.
  • E2E: a dataset of restaurant reviews, containing 42,061 training and 4693 test instances. We take sequence length 100.
B.2 Experiment settings
For images, we use $224\times 224$ resolution and patch 16 for vision transformers. After pre-training, an additional classifier head is inserted for each downstream task.
B.2.1 Figure 1
We use batch size 1000 and search over learning rate $\in[5e-5,1e-4,2e-4,5e-4]$.
  1. 1.
    CIFAR10, ViT-Base, random initialization, $\eta=5e-5$ without noise, $\eta=5e-4$ with noise, 10 epochs.
  2. 2.
    CIFAR100, ViT-Large, pretrained by timm library, $\eta=5e-5$ without noise, $\eta=5e-4$ with noise, 5 epochs.
  3. 3.
    CodeParrot, GPT2-Large, random initialization, $\eta=1e-4$ without clipping, $\eta=5e-4$ with clipping. This follows closely from https://huggingface.co/learn/nlp-course/chapter7/6.
  4. 4.
    E2E, GPT2-Large, pretrained by transformers library, $\eta=1e-4$ without noise, $\eta=5e-4$ with noise, 10 epochs.
B.2.2 Figure 4
$n=10^{6},\epsilon=1,\delta=1/n,S=10^{6}$ (1 epoch). We use the privacy accountants in Opacus [88]: RDP [59], GDP [26, 8], PRV [34]. The dashed lines depict $\sigma(1)^{2}/B$ for these privacy accountants.
B.2.3 Figure 5
For pre-training, $\sigma^{2}=0.25,{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}=1e2,\text{tr}({\mathbf{H}})/c^{2}=2e8,\text{tr}({\mathbf{H}}\mathbf{\Sigma})=2e4$; for fine-tuning, $\sigma^{2}=0.25,{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}=1e2,\text{tr}({\mathbf{H}})/c^{2}=2e6,\text{tr}({\mathbf{H}}\mathbf{\Sigma})=2e4$.
B.2.4 Figure 6
We train ViT-Base with $\eta=5e-4,B=500,\epsilon=2$. We use the Hutchinson method to compute Hessian-related trace, e.g. $\text{tr}({\mathbf{H}})=\mathbb{E}_{{\mathbf{v}}\in\mathbb{R}^{d}}[{\mathbf{v}}^{\top}{\mathbf{H}}{\mathbf{v}}]\leftarrow\frac{1}{k}\sum_{i=1}^{k}{\mathbf{v}}_{i}^{\top}{\mathbf{H}}{\mathbf{v}}_{i}$, where ${\mathbf{v}}\sim N(0,I_{d})$ and $k=100$. Note the Hessian-vector product ${\mathbf{H}}{\mathbf{v}}$ can be computed by one back-propagation of ${\mathbf{G}}({\mathbf{w}})^{\top}{\mathbf{v}}\in\mathbb{R}$ on ${\mathbf{w}}$.
B.2.5 Figure 7
We train GPT2-small with $\eta=2\times 10^{-4}$, $B=7680$, epochs $E_{\textup{pub}}=1$ (if not early stopped by patience $=2$), $E_{\textup{priv}}=3$ for mixed training, $E=4$ for other cases. In the figure, ‘PubRatio’ is the percentage of public data among all data.
B.2.6 Table 4
We empirically observe that ViT-Base models with/without fine-tuning on ImageNet1k.
E.g. vit_base_patch16_224_miil.in21k_ft_in1k v.s. vit_base_patch16_224_miil.in21k;
vit_base_patch16_224.augreg_in21k_ft_in1k v.s. vit_base_patch16_224.augreg_in21k are very similar in final accuracy, though those with fine-tuning converge faster at early epochs.
We train for 10 epochs, $B=1000$.
  • non-DP fine-tuning: full-parameter, $\eta=5e-5$ (except for SVHN and CIFAR10 under $\epsilon=2$, we use $\eta=2e-5$ instead).
  • DP fine-tuning: BiTFiT [12], $\eta=2e-3$ (except for SVHN and CIFAR10 under $\epsilon=2$, we use $\eta=5e-4$ instead). We observe that for our models, the accuracy may benefit from a smaller learning rate (1e-3) when $\epsilon\geq 8$.
B.2.7 Table 6
We train with full-parameter fine-tuning and $B=1000$. Our learning rate $\eta=5e-5$ ($\epsilon=8$) and $\eta=2e-5$ ($\epsilon=2$). $x$-shot means $x$ samples per class, e.g. 30-shot CIFAR100 means 3000 samples in total.
B.2.8 Table 6
We train with $\eta=5e-5$ ($\epsilon=8$) and $\eta=2e-5$ ($\epsilon=2$) for full-parameter fine-tuning and parameter-efficient fine-tuning. We note that with much heavier (public) pre-training, e.g. on JFT4B, DP NFnet-F3 and JFT4B can achieve even better accuracy [24].
B.3 Details in training stage switching
When switching from public pre-training to private continual pre-training, the DP gradient may have a different scale than the historical optimizer states and make the continual pre-training unstable. Therefore, we investigate different strategies to re-initialize the optimizer states during the switching. Throughout the training, we use AdamW, which have three states: the iteration index ($t$), the first-order momentum (${\mathbf{m}}$), and the second-order momentum (${\mathbf{v}}$).
In the experiment, we train a ViT-Base model from scratch on the CIFAR100 dataset. In the first three epochs, we use vanilla AdamW, then we switch to DP-AdamW and continue training for one epoch. During the switching, we fixed the learning rate and reset different states ($t,{\mathbf{v}},{\mathbf{m}}$) to zeros. The ablation study result is shown in Figure 8.
From the figure, we observe that re-initializing the first-order momentum ${\mathbf{m}}$ (R1) results in the best performance when switching from public to private training. For the other cases, we observe a performance drop when switching.
Figure 8: Ablation study of switching from non-DP to DP training with AdamW on CIFAR100 dataset. When switching ($T=1200$), we re-initialize different states in the AdamW optimizer in different linestyles. “R1”, “R2”, and “RS” indicate ${\mathbf{m}}$, ${\mathbf{v}}$ and $t$ are re-initialized, respectively. “N” indicates no re-initialization, and “Ref” is the reference behavior of continual training with non-DP AdamW.
B.4 Details for MIA
To conduct the membership inference attack (MIA) in Table 7, we employ a white-box attack with full access to model parameters and data, to evaluate the data protection by our DP pre-training. The attack has two major steps. Step I: construct MIA dataset. To construct the MIA dataset, we use the following procedures: 1) for each image in ImageNet-11k, we compute its output logits and loss, which serves as the feature of the MIA dataset; 2) we randomly select $50\%$ of the testing images and the same number of training images ($522,496$ samples) as the MIA test set, and the rest $50\%$ of the testing images and $10\%$ of the training images as MIA train set; 3) we label the training images as class “$1$” and testing images as class “00”. This creates the MIA dataset with 11k features and binary labels. Step II: evaluate MIA performance. After obtaining the MIA dataset, we fit a binary logistic regression with the MIA training set to classify whether an image belongs to the training set of ImageNet-11k (class “$1$”). We use the L-BFGS optimizer and class re-weighting to train the model for $50$ epochs. After training, we evaluate the performance of the training classification model on the MIA test dataset and compute the classification accuracy, precision, recall, F1-score, and AUROC of the classification task. The MIA attack procedures are illustrated in Figure 9.
Figure 9: The process of membership inference attack (MIA).
Appendix C Related works
Convergence analysis of DP training (the goals)
Recent works have attributed the slow DP convergence to the large number of model parameters [95, 48], the noise level added to gradients [84, 14, 86], and the per-sample gradient clipping [11, 21, 94]. Our analysis in (6) and (8) covers these factors as well as the choice of hyperparameters. We note that a key quantity of DP convergence is $\textup{tr}({\mathbf{H}})$, which implicitly covers the model dimension $d$ and is also analyzed in [54] to give a dimension-free generalization bound. Moreover, existing works mostly study the empirical convergence on the training set, in terms of the gradient norm $||{\mathbf{g}}_{t}||\to 0$, the parameter space ${\mathbf{w}}_{t}\to{\mathbf{w}}_{*}$, or the training loss $L_{t}\to L_{*}$ as $t\to\infty$. Our work focuses on the generalization performance, and pays particular attention on the data efficiency (as well as computation efficiency) that is rarely captured in the literature.
More on DP convergence analysis (the assumptions)
The convergence analysis of non-convex DP deep learning is very challenging and under-studied. There are roughly 3 routes to walk around the challenge: (1) Working on convex models instead (see [73, 23] and more in Sec 4.1.2 [66]), e.g. assuming some form of convexity or only optimizing the last layer (essentially a linear model). This route offers deep theoretical insights thanks to the simplified models, but generally fails to match the training dynamics nor the performance of deep learning. For instance, last-layer training may work reasonably well in computer vision but not so in language tasks, obtaining only 26 BLEU score for DP-GPT2 on E2E dataset compared to 63 BLEU score via DP full-parameter training. (2) Working on the continuous-time gradient flow, rather than the gradient descent in practice, in order to get rid of the per-sample gradient clipping or the Gaussian noise. This route essentially works with infinitely small learning rate, which falls short in the performance as SOTA DP models are trained with large learning rate [49]. (3) Assuming that the loss $L$ is Lipschitz continuous (i.e. the gradient is bounded) [3, 81, 82] or Lipscthiz smooth [21, 14, 86]. While both assumptions lead to some insights of the DP training behaviors, the Lipschitz constant is hardly calculable and time-dependent (maybe even diverging) in practice. Particularly, in the Lipschitz continuous setting, the per-sample gradient clipping is not used in DP deep learning, reducing the technical difficulty significantly but missing the gist of DP-SGD from a practitioner’s viewpoint.
In contrast, our work does not rely on these and many other assumptions in the literature, because we do not study the loss convergence along all iterations. We instead scrutinize the per-iteration loss improvement in (5). Put differently, we focus on the local behavior of DP optimizer rather than the global behavior.
Differentially private training
The literature of DP deep learning has predominantly focused on fine-tuning the pre-trained models (e.g. full-parameter [1, 49, 14] and PEFT [78, 89, 12, 58]). Importantly, DP fine-tuning (1) can be as performant as the standard non-DP fine-tuning consistently, (2) achieves better empirical performance with larger models [14, 9] and (3) necessarily relies on the public data pre-training. Our work is distinct from previous works since we focus on DP pre-training with large sample size, large number of iterations, and very limited public data. Interestingly, we notice that DP pre-training also enjoy these desirable characteristics of DP fine-tuning, e.g. requiring public data (only a small amount) in Section 4.1 and benefiting from scaling up the models (see Figure 4, Figure 5 and Table 9 in [90]). This empirical evidence of similar training dynamics between DP pre-training and DP fine-tuning, despite their difference in learning goals, is well-supported by our decelerator analysis in Section 3.
Continual pre-training
Continual pre-training is a training strategy that accumulates knowledge from a large amount of data, in order to improve its generalization ability to unseen domains and datasets [36, 28, 83, 43, 72]. It is different from fine-tuning which focuses on a task-specific and often smaller dataset. Therefore, the (continually) pre-trained models possess strong few-shot (includign zero-shot) ability but may not be competent in specific functionality such as conversation, and vice versa for fine-tuned models like ChatGPT, Alpaca [76], and Dolly [22].
Multiple DP pre-trained models have been developed [47, 24, 90], but these models either (1) suffer from low accuracy, e.g. 6.9% on ImageNet-1k without additional data by [47] and 32.4% by [24]; or (2) demand significantly more data and compute to match the non-DP pre-training, e.g. 100$\times$ in [90] when comparing DP ViP model to non-DP SimCLR; or (3) rely on uncommon tricks such as a huge batch size from 16-98K, that are not adopted in the standard deep learning community. Related to (3), we believe DP continual pre-training can further improve using the existing techniques from the non-DP continual training, including the replay mechanism and curriculum learning, and avoid catastrophic forgetting.
Hessian-based analysis in deep learning
Applying the second-order Taylor expansion of loss motivates the famous Newton’s method and is commonly adopted in deep learning [57, 97, 85, 56, 91], where the Hessian matrix is useful to analyzing the convergence and data efficiency (e.g. selecting the critical batch size [57]), even though it is infeasible to derive ${\mathbf{H}}\in\mathbb{R}^{d\times d}$ explicitly for large models. Our work follows the same path with a specific focus on DP related operations (i.e. the clipping and the noise). We use the Hutchinson method and Hessian-vector product to compute $\textup{tr}({\mathbf{H}})$, ${\mathbf{H}}{\mathbf{G}}$ and so on.
System design of DP training
To make DP deep learning broadly applicable, we believe it is necessary to not only evaluate DP algorithms on the utility, but also from a system design perspective. The design of a DP system, such as our DP continual pre-training, should resemble that of the standard non-DP system (see our extensive discussion in Remark 4.2). Such a design will be compatible to and benefit from new advances in the much larger non-DP literature, unifying DP and non-DP communities, instead of crafting techniques that are limited to DP learning only.
Data distribution shift
There may be some distribution shift between the public pre-training data and the private fine-tuning data or continual pre-training data. Empirically speaking, DP training can be robust to such distribution shift. For instance, DP fine-tuning has successfully transferred from ImageNet to CIFAR10, CIFAR100, SVHN, Food101, CelebA, FMNIST, GSTRB [14], from JFT to ImageNet [24, 58], and from Places365 to ImageNet [47]. Although this work does not address the distribution shift due to similarity between ImageNet-1k and ImageNet-11k, our analysis in Section 3, especially the formula of decelerator, still holds in this scenario. Therefore, we expect our DP continual pre-trainin to work withstanding the data distribution shift, as empirically demonstrated from Shaders to LAION [90].
Concurrent work
The concurrent work – ViP [90] – also studies the DP continual pre-training, thus is the closest strategy to ours. We elaborate the similarity and differences between two strategies.
Similarity
  1. 1.
    Both use ViT-Base as backbone and AdamW as optimizer.
  2. 2.
    Both use self-supervised learning (SSL) for public pre-training.
  3. 3.
    Both use $\epsilon=8$ for DP continual pre-training.
  4. 4.
    Both public datasets (Shaders v.s. ImageNet-1k) have $\approx 1$M images.
Differences
  1. 1.
    ViP uses SSL on DP continual pre-training. We use supervised learning.
  2. 2.
    ViP adds a decoder to ViT-Base (in total 99.0M parameters) during the pre-training. We replaces the classifier head of ViT-Base (94.4M parameters) for DP continual pre-training. Hence the model architecture is different.
  3. 3.
    ViP trains on Shaders (for 1.6B images) and LAION (for 0.6B images). We train on ImageNet-1k (for 0.3B images) and ImageNet-11k (for 0.4B images). Hence data distribution and quality is different, and our training requires about 1/3 the computation.
  4. 4.
    ViP’s private dataset (LAION) has 233M images. Ours (ImageNet-11k) has 12M images. Hence our training requires much smaller dataset (see Figure 3(a) [90]).
  5. 5.
    ViP has to use huge batch size 98K (see Figure 4(b) [90], where the loss diverges with batch size 8K). We use 4K.
  6. 6.
    ViP experiments with various model sizes, from ViT-Nano (18.6M) to Large (233M). We only use ViT-Base.
Appendix D Algorithm of DP continual pre-training
Algorithm 1 DP continual pre-training
1:switch_to_DP=False
2:for $t=1,2,...$ do
3:     Compute the loss $L_{t}$ by forward pass
4:     if switch_to_DP==False: then
5:         Compute public gradient $\mathbf{g}$
6:     else
7:         Compute private gradient $\mathbf{g}$
8:     end if
9:     Update $\mathbf{w}_{t+1}=\mathbf{w}_{t}-\eta\mathbf{g}$
10:     if $L_{t}>L_{t-1}$  then
11:         Set switch_to_DP=True
12:     end if
13:end for
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction accurately reflect the paper’s contributions.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We recognize the limitations of the second-order Taylor expansion and that we cannot explain the trace of Hessian’s evolution.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate ”Limitations” section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This paper has no theorem.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We give all the necessary information and will open-source our codebase.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Open access to code will be available after paper decision.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We give all details in the main text and in appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We believe the pattern is clear.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer ”Yes” if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [No]
  39. Justification: This work focuses on the accuracy, independent of the computer resources.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: It does.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: Not applicable.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: No such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite properly.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: Not applicable.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.